OpenStudy (anonymous):

Simplify 10 cubed root of 12 end root minus 7 fourth root of 12 end root minus 5 cubed root of 12 end root minus 2 fourth root of 12

OpenStudy (anonymous):

\[10\sqrt[3]{12} - 7\sqrt[4]{12} - 5\sqrt[3]{12} - 2\sqrt[4]{12} \] Like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well, just combine like terms, combine the cubed roots and then combine the 4th roots.

OpenStudy (anonymous):

okay thank you!